# Computational Sentence-level Metrics Predicting Human Sentence Comprehension

### v1

@@ Abstract
The majority of research in computational psycholinguistics has concentrated on the processing of words. This study introduces innovative methods for computing sentence-level metrics using multilingual large language models. The metrics developed—sentence surprisal and sentence relevance— are tested and compared to validate whether they can predict how humans comprehend sentences as a whole across languages. These metrics offer significant interpretability and achieve high accuracy in predicting human sentence reading speeds. Our results indicate that these computational sentence-level metrics are exceptionally effective at predicting and elucidating the processing difficulties encountered by readers in comprehending sentences as a whole across a variety of languages. Their impressive performance and generalization capabilities provide a promising avenue for future research in integrating LLMs and cognitive science.
@@ 1 Introduction
Although the related studies have claimed that some given computational metrics are predictive of sentence processing by humans, the fact is that these metrics can only predict how words are processed in a given context (such as word surprisal) (Hale, 2016; De Varda and Marelli, 2023; Gotlieb Wilcox et al., 2023). However, focusing solely on word processing may not truly mirror the real-world human experience of understanding sentences. In other words, we should focus on how humans comprehend entire sentences, rather than merely the individual words within them. It is necessary to broaden our understanding of sentence comprehension, placing human sentence processing within a larger context: discourse. Sentence-level metrics (such as sentence probability, sentence semantic relevance) are potentially useful in understanding entire sentence comprehension and processing by humans. For instance, the metric of contextual sentence-level relevance computed by Transformer-based language models could predict how Chinese natives comprehend sentences as a whole (Sun and Wang, 2022). A similar method can be applied in investigating sentence comprehension in other languages. In other words, we can enhance our analysis from word-level metrics to sentence-level metrics, exploring their ability to predict the processing of a sentence within a discourse-level context.
@@
Further, during naturalistic discourse reading, a number of factors may have an effect on sentence processing, including expectations concerning the next sentence based on the preceding context, and memory integrated with new input based on semantic relatedness with the context. Surprisal reflects the amount of information conveyed by a linguistic unit and its predictability, while semantic relevance represents the relatedness of a linguistic unit to the other units. Both factors could independently influence processing difficulty, but they may also interact, creating complex processing situations. The study aims to test hypotheses about the nature of these factors and understand their potentials in human sentence comprehension.
@@
With the advent of deep learning techniques and the availability of massive datasets, there is now an opportunity to develop and compute such sentence-level metrics. By using multi-lingual large language models (LLMs), we can estimate the probabilities of the next sentence (sentence surprisal) or compute semantic relevance among sentences (sentence relevance). These sentence-level metrics we proposed are expected to predict how humans comprehend and process sentence as a whole. Because of sentence length and human memory-stored limit, we consider a limited number of surrounding sentences as the context in calculating contextual information and further incorporate the contextual information in our metrics. The method is “attention-aware” approach because it is inspired by the attention mechanism in Transformer (Vaswani et al., 2017).
@@
We plan to apply these computational metrics to test their predictability in a number of languages rather than in English. Fortunately, such datasets and computational tools are available. For instance, existing multi-lingual LLMs are highly capable of processing and understanding text in multiple languages, further helping compute sentence-level metrics we propose. There are also multilingual databases on language comprehension and processing. For instance, the Multilingual Eye-movement Corpus (MECO) is a collection of eye-tracking data that has been collected from participants reading texts in 13 languages (Siegelman et al., 2022). This is an ideal testing dataset for our research. The present study aims to predict sentence comprehension by using sentence-level computational metrics. By employing these metrics, we expect to obtain more accurate predictions of how sentences are processed and comprehended holistically. To evaluate the generalizability of our approach to various languages, we will undertake a multi-lingual investigation.
@@ 2 Related Work
Word surprisal, the negative logarithm of a word’s probability in left context, is effectively used in testing expectation-based theories and predicting language processing (Hale, 2016). This metric signifies that predictable words require less cognitive effort to process. However, it remains an open question: which context aspects - lexical, syntactic, semantic, or conceptual - facilitate prediction or integration. As the language models developed so quickly, the methods for calculating word probabilities have been revolutionized in the transition from n-gram to Transformer (Wolf et al., 2020). For instance, recently deep-learning models have dominated many NLP tasks and LLMs have been used as models for calculating the processing difficulty (Goodkind and Bicknell, 2018; Wilcox et al., 2020; Schrimpf et al., 2021). The recent LLMs are effective and convince to compute the probabilities of the next sentence in context. Moreover, “surprisal” is characterized as the negative logarithm of the probability of an event, and surprisal can be applied across different linguistic levels, including sentences, discourse, or phonemes (Venhuizen et al., 2019; Pimentel et al., 2021; Gwilliams and Davis, 2022). Thus, provided we have sufficient computational resources, we can accurately calculate sentence-level surprisals for mutiple languages.
Additionally, some effective methods have been proposed for computing semantic relevance among words within a sentence. The new approach involves integrating contextual information using the attention-aware method, resulting in more powerful contextual semantic relevance (Sun et al., 2023a; Sun et al., 2023b). However, we believe that the new method can be extended and modified to process entire sentences rather than merely words. By employing an attention-aware approach, we can calculate the semantic relevance among sentences, enabling us to measure the extent to which a target sentence is semantically connected to its neighboring sentences within the discourse.
@@ 3.1 Testing datasets
Multilingual databases facilitate cost-effective and reliable cognitive and neural language testing. The Multilingual Eye-tracking Corpus (MECO) is particularly useful due to its diverse range of 13 languages and eye-tracking data indicating cognitive effort in processing (Siegelman et al., 2022). Each language’s participants read 12 encyclopedia entries of 2000 tokens, resulting in 36000 tokens in total. These entries were similarly complex across languages and generated about 70,000 to 80,000 eye-tracking data points per language. MECO was chosen for its language diversity, large native speaker counts, and ample eye-tracking data. Reading speed (or rate) has been extensively explored, typically employing text or individual sentence as the unit of interest (Miller and Coleman, 1971; Carver, 1976; Biancarosa, 2005; Brysbaert, 2019; Siegelman et al., 2022). This study focuses on sentence reading speed (= the number of words in this sentence / total fixation duration of a sentence) - word number per second. When the reading speed is low, readers use more time to process this sentence. In contrast, a higher sentence reading speed indicates that less time is used to process every word for this sentence. While reading speed is not a direct oculomotor measure, it is influenced by oculomotor factors such as fixation duration, saccade durations, and the tendency to skip words. Reading speed is interconnected with these oculomotor measures, as they collectively contribute to how quickly and efficiently one can read and understand a sentence or text. In this sense, reading speed is a broader measure of language comprehension and processing. The reading speed is therefore considered as the overall difficulty or speed when readers comprehend a sentence/text as a whole.
@@ 3.2 Computing sentence surprisal
We employed two multilingual LLMs: m-BERT and mGPT to compute sentence surprsial. Sentence surprisal is the negative logarithm of next sentence probability ($-log(p(sentence|left\quad context))$), which is similar to word surprisal. Multi-lingual BERT (m-BERT) (Devlin et al., 2018) can be employed to compute word-level or sentence-level surprisal for various languages. BERT, a masked language model, has good performance in a number of NLP tasks (Salazar et al., 2019; Kalyan et al., 2021), and can be used to estimate next word probabilities. We used the state-of-the-art BERT model (i.e., multilingual-bert-uncased) because it can be consistently applied in different languages in order to compute surprisal. However, it seems a little difficult to use it to compute sentence probability immediately using m-BERT. Fortunately, we can take some strategies to allow m-BERT to approximate next sentence probabilities. For instance, we gauged the joint probability of an entire sentence conditioned on its preceding context. This is achieved by using the chain rule of probability, breaking the sentence into individual tokens, and computing the probability of each token given all preceding tokens (and the context).
@@
To compute sentence surprisal, we also proposed the other methods to compute sentence surprisal. The three methods are summarized as follows: the first one is to use chain rule (CR), which was mentioned above. Sentence surprisal is calculated by tokenizing a sentence into individual elements and computing the joint probability of the sentence given a context. This is done by multiplying the conditional probabilities of each token given its preceding tokens and the context, applying the chain rule of probability to language sequences. The method can be computed based on either m-BERT or mGPT. The second method is to use “next sentence prediction” (NSP mechanism in m-BERT to compute sentence surprsial. Third, sentence surprisal is also quantified by computing the Negative Log-Likelihood (NLL) of the sentence when conditioned on its context. NLL is used as a loss function to measure how well the model’s predictions align with the actual data. For a given sequence of tokens, NLL is a measure of how surprised the model is by the actual sequence. The details on the three methods are seen in Appendix A.
Overall, we employed three distinct methods for computing sentence surprisal, utilizing two multilingual LLMs: m-BERT and mGPT. “ Sentence surprisal” quantifies the level of unpredictability or information associated with encountering a given sentence. The CR and the NLL approach were applied to both LLMs for this purpose. In contrast, the NSP method was exclusively implemented on m-BERT. The methods and models are shown in the Panel A of Fig. 1.
@@ 3.3 Computing attention-aware sentence relevance
This section elaborates on the computation of sentence semantic relevance. Consider a window comprising four sentences, as depicted in Panel B of Fig. 1, where the objective is to evaluate the semantic relationship of the target sentence with the adjacent three sentences. The computation unfolds in two primary steps. Initially, embeddings for each sentence within the window are generated using either m-BERT or mGPT. Subsequently, the application of cosine similarity to the sentence embeddings derived from m-BERT or mGPT serves as a conventional approach to ascertain the semantic similarity between any pair of sentences, as detailed in Appendix B. Nonetheless, this step is to obtain the relevance of two sentences in this window. Our objective, however, is to compute the semantic relatedness between a target sentence and its contextual sentences, which is another sentence-level metric.
Upon calculating the semantic similarity values for any pair of sentences within this window, it is necessary to employ an “attention-aware” approach to process these values across all sentences. The “attention-aware” approach has been successfully applied in computing word-level metrics for predicting reading (Sun et al., 2023a; Sun et al., 2023b). Specifically, we look at the preceding two sentences and the following sentence. The similarity score for any two sentences is computed by taking the dot product between the first embedding of the concatenated input and the second embedding of the concatenated input. Subsequently, we derived four cosine similarity values for these sentences. Each value is then weighted differently based on its distance from the target sentence. It is formalized as (1).
@@
Here “W” represents weights according to the positional distance between the target sentence (“s”) and its contextual sentences (“c”). The range of weights is between 1/2 and 1/3 (i.e the neighboring sentence is multiplied by 1/2, but the non-neighboring one is by 1/3.). As the distance from the target sentence is large, the weight is given with smaller one. The weights used for attention-aware metrics mimic human forgetting mechanism, a representation of memory retention decline over time, where retained information halves after each day within a span of several days (Loftus, 1985), shown in Fig. 3. Unlike the attention weights in transformers, which are computed via neural networks and are not easily interpretable. Each cosine similarity value is multiplied by its corresponding weight, and the results are then aggregated to produce a single value. This value signifies the semantic relevance of the target sentence within its context(i.e., “sentence relevance”). Leveraging various sentence embeddings created by m-BERT and mGPT, we are able to derive two distinct metrics of “sentence relevance” for the identical target sentence within its specific context. The computation is shown as Panel B of Fig. 1.
The method proposed for computing attention-aware metrics works as a memory agent (see Appendix B). This approach considers both preceding and following sentences as potential sources of contextual information. The relationship between these contextual sentences and the target sentence is then weighted based on their positional distance and aggregated to create a comprehensive measure of semantic relevance within the discourse. This approach is highly explainable from linguistic and cognitive perspectives. Table 1 provides an overview of the metrics in our analysis.
Table 1: The models and statistical analysis used in this study
@@ 3.4 Statistical method and model comparison
We employed Generalized Additive Mixed Models (GAMMs) (Wood, 2017) to investigate the predictive power of certain measures on eye-movements during reading. The performance of GAMM models was evaluated using difference in AIC (Akaike’s Information Criterion) (i.e.,$\Delta$AIC ) between a base GAMM and a full GAMM served as a measure to assess the effectiveness of a metric. A smaller $\Delta$AIC value indicates that the measure we are interested in estimates provide more accurate predictions of the response variable compared to the baseline model.The difference in AIC between a base GAMM and a full GAMM served as a measure to assess the effectiveness of a computational measure. A smaller (negative) $\Delta$AIC
value indicates that the measure we are interested in estimates provide more accurate predictions of fixation duration compared to the baseline model.
Comparing GAMM models, a larger model will have a higher likelihood due to more parameters. But, comparing loglikelihoods (or log-likelihood (LogLik)) directly is not suitable for differently sized models. Both AIC and Bayesian Information Criterion (BIC) address this by penalizing model complexity. However, AIC is preferred for prediction quality, while BIC prioritizes parsimony, often favoring smaller models. Previous studies (Wilcox et al. (2020), Oh and Schuler (2023)) used $\Delta$LogLik to analyze the predictive power of surprisal computed various language models. However, LogLik may not be a good standard. Moreover, including random variables (e.g., participants) and significant predictors like “word length” and “word frequency” is crucial for accurate regression model evaluation. Failure to do so may result in suboptimal fittings for studying reading time. Without optimization, there is a risk of encountering misleading issues that can undermine the validity of the results. The regression models in these studies (Wilcox et al., 2020; Oh and Schuler (2023)) may not be optimal because of lack of random variables or control predictors. The present study included these control predictors (word length/frequency) and random variable (e.g.,“languages”, “participants” ) in the GAMMs. Furthermore, the excessive imposition of heavy penalties can lead to overfitting in GAMM models. We have avoided this practice. More details are seen in Appendix C.

### v2

# Computational Sentence-level Metrics for Predicting Comprehension of Entire Sentence by Humans

@@ Abstract
The majority of research in computational psycholinguistics has concentrated on the processing of words. This study introduces innovative methods for computing sentence-level metrics using multilingual large language models. The metrics developed—sentence surprisal and sentence relevance— are tested and compared to validate whether they can predict how humans comprehend sentences as a whole across languages. These metrics offer significant interpretability and achieve high accuracy in predicting human sentence reading speed. Our results indicate that these computational sentence-level metrics are exceptionally effective at predicting and elucidating the processing difficulties encountered by readers in comprehending sentences as a whole across a variety of languages. Their impressive performance and generalization capabilities provide a promising avenue for future research in integrating LLMs and cognitive science.
@@ 1 Introduction
Although the related studies have claimed that some given computational metrics are predictive of sentence processing by humans, the fact is that these metrics can only predict how words are processed in a given context (such as word surprisal) (Hale, 2016; De Varda and Marelli, 2023; Gotlieb Wilcox et al., 2023). Focusing solely on word processing may not truly mirror the real-world human experience of comprehending sentences. In other words, we should also explore how humans comprehend entire sentences, rather than merely the individual words within them. It is necessary to broaden our understanding of sentence comprehension, placing human sentence processing within a larger context: discourse. Computational sentence-level metrics (such as sentence probability, sentence semantic relevance) are potentially useful in understanding entire sentence comprehension and processing by humans. For instance, the metric of contextual sentence-level relevance computed by Transformer-based language models could predict how Chinese natives comprehend sentences as a whole (Sun and Wang, 2022). A similar method can be applied in investigating sentence comprehension in other languages. In short, we can enhance our analysis from word-level metrics to sentence-level metrics, exploring their predictability for the processing of a sentence within a discourse-level context.
@@
During naturalistic discourse reading, a number of factors may have an effect on sentence processing, including expectations concerning the next sentence based on the preceding context, and memory integrated with new input based on semantic relatedness with the context. Surprisal reflects the amount of information conveyed by a linguistic unit and its predictability, while semantic relevance represents the relatedness of a linguistic unit to the other units. Both factors could independently influence processing difficulty, but they may also interact, creating complex processing situations. The study aims to compute these metrics at the senentece level and test hypotheses about their nature and understand their potentials in human sentence comprehension.
@@
With the advent of deep learning techniques and the availability of massive datasets, there is now an opportunity to develop and compute such sentence-level metrics. By using multi-lingual large language models (LLMs), we can estimate the probabilities of the next sentence (sentence surprisal) or compute semantic relevance among sentences (sentence relevance). These sentence-level metrics we proposed are expected to predict how humans comprehend and process sentence as a whole. Because of sentence length and human memory-stored limit, we consider a limited number of surrounding sentences as the context in calculating contextual information and further incorporate the contextual information in our metrics. The method is “attention-aware” approach because it works as effectively as the attention mechanism in Transformer (Vaswani et al., 2017) in processing contextual information.
@@
We plan to apply these computational metrics to test their predictability in a number of languages rather than in English. Fortunately, such datasets and computational tools are available. For instance, existing multi-lingual LLMs are highly capable of processing and understanding text in multiple languages, further helping compute sentence-level metrics we propose. There are also multilingual databases on language comprehension and processing. For instance, the Multilingual Eye-movement Corpus (MECO) is a collection of eye-tracking data that has been collected from participants reading texts in 13 languages (Siegelman et al., 2022). This is an ideal testing dataset for our research. The present study aims to compute sentence-level computational metrics and apply them to predicting sentence reading speed. By employing these metrics, we expect to obtain more accurate predictions of how sentences are processed and comprehended holistically. To evaluate the generalizability of our approach to various languages, we will undertake a multi-lingual investigation.
@@ 2 Related Work
Word surprisal, the negative logarithm of a word’s probability in left context, is effectively used in testing expectation-based theories and predicting language processing (Hale, 2016). This metric signifies that predictable words require less cognitive effort to process. However, it remains an open question: which context aspects - lexical, syntactic, semantic, or conceptual - facilitate prediction or integration. While word surprisal accounts for the influence of the left context, this context is confined to a single sentence. Word surprisal has primarily been used to predict individual word processing times, such as the duration of reading a word. However, it falls short in predicting sentence-level measures, such as overall reading speed. Therefore, our goal is to develop metrics that go beyond within-sentence analysis to accurately predict reading speed. As the language models developed so quickly, the methods for calculating word probabilities have been revolutionized in the transition from n-gram to Transformer (Wolf et al., 2020). For instance, recently deep-learning models have dominated NLP tasks and LLMs have been used as models for calculating the processing difficulty (Goodkind and Bicknell, 2018; Wilcox et al., 2020; Schrimpf et al., 2021). The recent LLMs are effective to compute the probabilities of the next sentence in context. Moreover, “surprisal” is characterized as the negative logarithm of the probability of an event, and surprisal can be applied across different linguistic levels, including sentences, discourse, or phonemes (Venhuizen et al., 2019; Pimentel et al., 2021; Gwilliams and Davis, 2022). Provided we have sufficient computational resources, we can accurately calculate sentence-level surprisals for mutiple languages.
Moreover, some effective methods have been proposed for computing semantic relevance among words within a sentence. The new approach involves integrating contextual information using the attention-aware method, resulting in more powerful contextual semantic relevance (Sun et al., 2023a; Sun et al., 2023b). However, we believe that the similar methods can be extended and modified to process entire sentences rather than merely words. By employing an “attention-aware” approach, we can calculate the semantic relevance among sentences, enabling us to measure the extent to which a target sentence is semantically connected to its neighboring sentences within the discourse.
@@ 3.1 Testing datasets
Multilingual databases facilitate cost-effective and reliable cognitive and neural language testing. The Multilingual Eye-tracking Corpus (MECO) is particularly useful due to its diverse range of 13 languages and eye-tracking data indicating cognitive effort in processing (Siegelman et al., 2022). Each language’s participants read 12 encyclopedia entries of 2000 tokens, resulting in 36000 tokens in total. These entries were similarly complex across languages and generated about 70,000 to 80,000 eye-tracking data points per language. MECO was chosen for its language diversity, large native speaker counts, and ample eye-tracking data. Reading speed (or rate) has been extensively explored, typically employing text or individual sentence as the unit of interest (Miller and Coleman, 1971; Carver, 1976; Biancarosa, 2005; Brysbaert, 2019; Siegelman et al., 2022).
This study focuses on sentence reading speed (= the number of words in this sentence / total fixation duration of a sentence) - word number per second (for a sentence or text). When the reading speed is low, readers use more time to process this sentence. In contrast, a higher sentence reading speed indicates that less time is used to process every word for this sentence. While reading speed is not a direct oculomotor measure, it is influenced by oculomotor factors such as fixation duration, saccade durations, and the tendency to skip words. Reading speed is interconnected with these oculomotor measures, as they collectively contribute to how quickly and efficiently one can read and understand a sentence or text. Fixation duration for word merely represents the difficulty for an individual word. In contrast, the reading speed is considered as the overall difficulty when readers comprehend a sentence/text as a whole. While reading speed measures the number of words processed per minute or second across sentences, word reading duration assesses the processing difficulty of a single word. As these measures are fundamentally different, word surprisal predicts the processing of individual words and does not extend to sentence-level processing (More are seen in Appendix A). We need to figure out practical sentence-level metrics to estimate reading speed.
@@ 3.2 Computing sentence surprisal
We employed two multilingual LLMs: m-BERT and mGPT to compute sentence surprsial. Sentence surprisal is the negative logarithm of next sentence probability ($-log(p(sentence|left\quad context))$), which is similar to word surprisal. Multi-lingual BERT (m-BERT) (Devlin et al., 2018) can be employed to compute word-level or sentence-level surprisal for various languages. BERT, a masked language model, has good performance in a number of NLP tasks (Salazar et al., 2019; Kalyan et al., 2021), and can be used to estimate next word probabilities. We used the state-of-the-art BERT model (i.e., multilingual-bert-uncased) because it can be consistently applied in different languages in order to compute word surprisal. However, it seems a little difficult to use it to compute sentence probability immediately using m-BERT. Fortunately, we can take some strategies to allow m-BERT to approximate next sentence probabilities. For instance, we gauged the joint probability of an entire sentence conditioned on its preceding context. This is achieved by using the chain rule of probability, breaking the sentence into individual tokens, and computing the probability of each token given all preceding tokens (and the context).
@@
To compute sentence surprisal, we also proposed the other methods to compute sentence surprisal. The three methods are summarized as follows: the first one is to use chain rule (CR), which was mentioned above. Sentence surprisal is calculated by tokenizing a sentence into individual elements and computing the joint probability of the sentence given a context. This is done by multiplying the conditional probabilities of each token given its preceding tokens and the context, applying the chain rule of probability to language sequences. The method can be computed based on either m-BERT or mGPT. The second method is to use “next sentence prediction” (NSP) mechanism in m-BERT to compute sentence surprsial. Third, sentence surprisal is also quantified by computing the Negative Log-Likelihood (NLL) of the sentence when conditioned on its context. NLL is used as a loss function to measure how well the model’s predictions align with the actual data. For a given sequence of tokens, NLL is a measure of how surprised the model is by the actual sequence. The details on the three methods are seen in Appendices B & C.
Overall, we employed three methods for computing sentence surprisal, utilizing two multilingual LLMs: m-BERT and mGPT. “Sentence surprisal” quantifies the level of unpredictability or information associated with encountering a given sentence. The CR and the NLL approach were applied to both LLMs for this purpose. In contrast, the NSP method was exclusively implemented on m-BERT. The methods and models are shown in the Panel A of Fig. 1.
@@ 3.3 Computing attention-aware sentence relevance
This section elaborates on the computation of sentence semantic relevance. Consider a window comprising four sentences, as depicted in Panel B of Fig. 1, where the objective is to evaluate the semantic relationship of the target sentence with the adjacent three sentences within a window. The computation unfolds in two primary steps. Initially, embeddings for each sentence within the window are generated using either m-BERT or mGPT. Subsequently, the application of cosine similarity to the sentence embeddings derived from m-BERT or mGPT serves as a conventional approach to ascertain the semantic similarity between any pair of sentences, as detailed in Appendix D. Nonetheless, this step is to obtain the relevance of two sentences in this window. Our objective, however, is to compute the semantic relatedness between a target sentence and its contextual sentences, which is another sentence-level metric.
Upon calculating the semantic similarity values for any pair of sentences within this window, we employ an “attention-aware” approach to process the four similarity values. The “attention-aware” approach has been successfully applied in computing word-level metrics for predicting reading (Sun et al., 2023a; Sun et al., 2023b). Specifically, we look at the preceding two sentences and the following sentence. Subsequently, we obtain four cosine similarity values for these sentences. Each value is then weighted differently based on its distance from the target sentence. It is formalized as (1).
@@
Here “W” represents weights according to the positional distance between the target sentence (“s”) and its contextual sentences (“c”). The range of weights is between 1/2 and 1/3 (i.e the neighboring sentence is multiplied by 1/2, but the non-neighboring one is by 1/3.). As the distance from the target sentence is large, the weight is given with smaller one. The weights used for attention-aware metrics mimic human forgetting mechanism, a representation of memory retention decline over time (e.g., the mean decline rate is 1/3 for three days), where retained information halves after each day within a span of several days (Loftus, 1985), shown in Fig. 4 (see Appendix D). Unlike the attention weights in transformers, which are computed via neural networks and are not easily interpretable. Each similarity value is multiplied by its corresponding weight, and the results are then aggregated to produce a single value. This value signifies the semantic relevance of the target sentence within its context(i.e., “sentence relevance”). Leveraging various sentence embeddings created by m-BERT and mGPT, we are able to derive two distinct metrics of “sentence relevance” for the identical target sentence within its specific context. The computation is shown as Panel B of Fig. 1.
The method proposed for computing attention-aware metrics was not invloved in introducing attention layers in Transformer. The term “attention-aware” sounds similar to attention because the method works as well as attention in incorporating contextual information. The “attention-aware” approach can work as a memory agent (see Appendix D). This approach considers both preceding and following sentences as potential sources of contextual information. The relationship between these contextual sentences and the target sentence is then weighted based on their positional distance and aggregated to create a comprehensive measure of semantic relevance within the discourse. This approach is highly explainable from linguistic and cognitive perspectives. Table 1 provides an overview of the metrics in our analysis.
Table 1: The models and statistical analysis
@@ 3.4 Statistical method and model comparison
We employed Generalized Additive Mixed Models (GAMMs) (Wood, 2017) to investigate the predictive power of sentence reading speed. The present study included these control predictors (mean word length/frequency for a sentence) and random variable (e.g.,“languages”, “participants” ). Including these variables is crucial for achieving optimal GAMM fitting.
The performance of GAMM models was evaluated using difference in AIC (Akaike’s Information Criterion) (i.e.,$\Delta$AIC ) between a base GAMM and a full GAMM served as a measure to assess the effectiveness of a metric. A smaller $\Delta$AIC value indicates that the measure we are interested in estimates provide more accurate predictions of the response variable compared to the baseline model. The difference in AIC between a base GAMM and a full GAMM served as a measure to assess the effectiveness of a computational measure. A smaller (negative) $\Delta$AIC value indicates that the measure we are interested in estimates provide more accurate predictions of fixation duration compared to the baseline model. The more details on statistical methods can be seen in Appendix E.